### v1

@@ Abstract
Shared effort games model people investing resources in public projects, where the share of the generated values is predefined. In linear $\theta$ sharing (effort) games, a project’s value is linear in the total contribution, while the threshold $\theta$ for effort defines which contributors win and receive their (equal) share. Thresholds between 0 and 1 model games such as paper co-authorship and shared assignments, where a minimum positive contribution is required for sharing in the value. We constructively characterise the conditions for the existence of a pure equilibrium for $\theta\in\{0,1\}$, and for two-player games with a general threshold, and find the prices of anarchy and stability. For more players, we also provide existence and efficiency results, and use generalised fictitious play simulations to show when a pure equilibrium exists and what its efficiency is. We also prove mixed equilibria always exist and bound their efficiency. This facilitates reaching socially efficient equilibria.
@@
Shared effort games, Contribution threshold, Equilibrium, Price of anarchy, Price of stability, fictitious play
journal: Games and Economic Behavior
@@ 1 Introduction
Many real-world situations include a group of players investing resources across multiple projects. The revenues of these projects are typically divided based on the investments. Examples of such situations include contributions to online communities Harper et al. (2007), Wikipedia Forte and Bruckman (2005), political campaigns Siegel (2009), paper co-authorship Kleinberg and Oren (2011), or social exchange networks Kleinberg and Tardos (2008). In the formal analyses of these type of games, it is often assumed that the obtained revenue from such projects is linear in the total contribution and is shared equally Roberson (2006); Kleinberg and Oren (2011); Bachrach et al. (2011). However, in most of the above examples, revenue is only shared among those who contribute at least a certain amount of effort. In this work we model this as a threshold on a player’s contribution. In particular, we analyse the situation when the considered threshold is relative to the investments of the other players.
An example with a relative threshold of $1$ is an all pay-auction. Another example is the Colonel Blotto game (see e.g. Roberson (2006)), where there are several battlefields over which each of the $2$ players distributes her soldiers, and the winner at each battlefield is the one who has put more people there. A player’s payoff is the number of projects (battlefields) where she has won, the value of each project being fixed. Thus, when our model is reduced to $2$ players with a fixed value of each project, we obtain the Colonel Blotto competition, whereas we typically assume a project has a value that increases as a function of the total contribution, modelling constructive collaboration, rather than countenance. These examples are “highly thresholded” because only the player whose effort per project is maximum collects the complete revenue. The following example with a smaller threshold is used later to further illustrate the model.
@@
It is important to understand and predict the stable contribution strategies, and suggest the efficient ones in situations where players invest resources in several projects and share the generated values. The classical stability model is the Nash equilibria, and the efficiency model of stable situations is the ratio of the least total utility of the players in an equilibrium to the optimum, called the price of anarchy ($\operatorname{PoA}$Koutsoupias and Papadimitriou (2009), and the ratio of the largest total utility in equilibrium to the optimal total utility, called the price of stability ($\operatorname{PoS}$Schulz and Moses (2003); Anshelevich et al. (2004). If the price of anarchy is close to $1$, then all equilibria are good, and we may follow any equilibrium profile. If the price of anarchy is low, while the price of stability is high, then we have to regulate the play by suggesting the efficient equilibria, while if even the price of stability is low, the only way to make the play socially efficient is changing the game through subsidizing, etc.
This price of anarchy for shared effort games was bounded in Bachrach et al. (2011), but assuming that a player obtains at least $1/k$th of her marginal contribution, which fails to hold in a positively thresholded model. Threshold was introduced in Polevoy et al. (2014), and pure equilibria for two players with close budgets and projects with linear utility functions were characterised. There is no analysis of the existence of Nash Equilibria (NE) and their efficiency in general linear shared effort games and nothing about mixed Nash equilibria is known. This paper aims to fill this gap by the following contributions: [^4] Relatively to Polevoy et al. (2014), we extend the theory also for budgets not within a threshold factor from each other, characterise the non-suppressed equilibria for threshold $1$, substantially extend the simulations, and prove the existence of an NE in the mixed case, answering this natural question.
@@
Analysing the cases of 00 threshold and $1$ threshold (Theorem 2), sufficiency results on existence of pure NE and efficiency bounds, and exploring this domain through simulation of fictitious play to find pure NE and their efficiency in 2-project multi-player games. This involves designing an $O(n\log n)$ best response algorithm for $2$-project multi-player games and proving that for more projects, best-responding is NP-hard, even for the same project functions and $2$ players.
@@
a proof of existence of mixed NE (it is an infinite game, so Nash’s Theorem Nash (1951) does not apply) and extending efficiency bounds to mixed equilibria.
After defining shared effort games in the next section, we first concentrate on pure strategies. We theoretically treat the existence and efficiency of NE for thresholds 00 and $1$ and then for games with two players in Section 3. Afterwards, we study games with any number of players. Aiming to further investigate existence and efficiency of pure NE for more than $2$ players, we employ fictitious play simulations in Section 4. Concretely, we define an Infinite-Strategy Fictitious Play and simulate it till and if it finds an equilibrium within some time. We check whether we have found an NE and if that is the case, what its efficiency is. Finally, we treat the mixed extension in Section 5, proving the existence of a mixed NE and showing some efficiency bounds of the pure case generalise to the mixed case. We conclude and discuss the future work in Section 7. When a proof is mostly technical, we provide its intuition in the body of the paper and defer the proof to A.
@@ 2 Model
To model investing effort in shared projects, we define shared effort games, extending the model by Bachrach et al. (2011) with thresholds. The games consist of players who contribute to project, and share the value the projects generates. Henceforth, we use the general term “value” instead of the monetary “revenue”. Formally, players $N=\left\{1,\ldots,n\right\}$ contribute to projects $\Omega$. Each player $i\in N$ can contribute to projects in $\Omega_{i}$, where $\emptyset\subsetneq\Omega_{i}\subseteq\Omega$; the contribution of player $i$ to project $\omega\in\Omega_{i}$ is denoted by $x^{i}_{\omega}\in{\mathbb{R}_{+}}$; for any $\omega\notin\Omega_{i}$, we write $x^{i}_{\omega}=0$. Each player $i$ has a budget $B_{i}>0$, and the strategy space of player $i$ (i.e., the set of her possible actions) is [^5] We denote “defined as” by =$\stackrel{{\scriptstyle\rm\Delta}}{{=}}$. $S^{i}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{x^{i}=(x^{i}_{\omega})_{\omega\in\Omega}\in\mathbb{R}_{+}^{\left|\Omega\right|}\mid\sum_{\omega\in\Omega_{i}}{x^{i}_{\omega}}\leq B_{i},\omega\notin\Omega_{i}\Rightarrow x^{i}_{\omega}=0\right\}$. Assume w.l.o.g. that $B_{n}\geq\ldots\geq B_{2}\geq B_{1}>0$. Being an $\left|\Omega\right|-1$-dimensional simplex, this set is compact and convex. Denote the vector of all the contributions by $x=(x^{i}_{\omega})^{i\in N}_{\omega\in\Omega}$ and the strategies of all the players except $i$ by $x^{-i}$.
To define the utilities, each project $\omega\in\Omega$ is associated with its linear (allows concentrating on the influence of interaction while modelling many “projects”, like research avenues, Wikipedia or environmental activism) project function, which determines its value, based on the total contribution $x_{\omega}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{i\in N}{(x_{\omega}^{i})}$ that the project receives; formally, $P_{\omega}(x_{\omega})\colon{\mathbb{R}_{+}}\to{\mathbb{R}_{+}}$, and $P_{j}(y)=\alpha_{j}y$. W.l.o.g., $\alpha_{m}\geq\alpha_{m-1}\geq\ldots\geq\alpha_{1}>0$. We denote the number of projects with the largest coefficient of project functions by $k\in\mathbb{N}$, i.e. $\alpha_{m}=\alpha_{m-1}=\ldots=\alpha_{m-k+1}>\alpha_{m-k}\geq\alpha_{m-k-1}\geq\ldots\geq\alpha_{1}$. We call those $k$ projects steep. The project’s value is distributed among the players in $N_{\omega}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{i\in N\mid\omega\in\Omega_{i}\right\}$ according to the following rule. From each project $\omega\in\Omega_{i}$, each player $i$ that contributes enough gets an equal share, denoted $\phi_{\omega}^{i}(x_{\omega})\colon{\mathbb{R}_{+}}\to{\mathbb{R}_{+}}$. Such sharing is relevant to many applications where a minimum contribution is required to share the value, such as paper co-authorship and homeworks, and we study predominantly such games. For any $\theta\in[0,1]$, define the players who get a share as those who bid at least a $\theta$ fraction of the maximum bid size to $\omega$,
@@
[…]
@@
$u^{i}(x)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{\omega\in\Omega_{i}}{\phi^{i}_{\omega}(x_{\omega})}.$
@@ 3 Pure Nash Equilibrium
We begin with general results on equilibria in shared effort games. Then, we completely characterise the existence and efficiency of Nash equilibria for two players, following with existence and efficiency results for any number of players. We study pure NE in this section, which are natural to play, since a pure shared effort game already has uncountably infinite strategy spaces and non-continuous utilities.
@@
[…]
Proof. The strategy space of player $i$ is $S^{i}$, and denote $S\stackrel{{\scriptstyle\rm\Delta}}{{=}}S^{1}\times\ldots\times S^{n}$. Define $P\colon S\to\mathbb{R}$ by $P(x)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{\omega\in\Omega}{\frac{P_{\omega}(x_{\omega})}{\left|N_{\omega}\right|}}$. This is a potential function, because it is equal to the utility of any player up to the terms from the projects out of $\Omega_{i}$, which are constant with respect to the player’s actions, and therefore, when player $i$ changes her strategy, her utility changes exactly as the potential does.
The game possesses a pure NE, whenever the potential function attains its maximum. As the linear project functions are continuous and the spaces are non-empty compact, being simplexes, so the potential function always attains its maximum (see Lemma $4.3$ in Monderer and Shapley (1996b)).
@@
[…]
@@
[…]
@@
We are finally ready to characterise the existence of an NE for two players.
Remark 1.
From now on in this section we assume that $\Omega_{i}$ includes all the steep projects, for each player $i\in N$, that is $\left\{m,m-1,\ldots,m-k+1\right\}\subseteq\Omega_{i},\forall i\in N$.
@@
In order to facilitate decisions, it is important to analyse the efficiency of the various Nash Equilibria. We aim to find the famous price of anarchy ($\operatorname{PoA}$), which is the ratio of a worst NE’s efficiency to the optimum possible one, and the price of stability ($\operatorname{PoS}$), which is the ratio of a best NE’s efficiency to the optimum possible one. We first completely handle two players.
@@
      $B_{1}<\frac{\theta B_{2}}{k}$ and $\frac{\alpha_{m-k}}{\alpha_{m}}\leq\min\left\{\frac{1}{1+\theta},\frac{2\theta}{1+\theta}\right\}$. Then, there exists a pure strategy NE and the following holds. $\operatorname{PoS}=\frac{\alpha_{m}B_{2}+\alpha_{m-k}B_{1}}{\alpha_{m}(B_{1}+B_{2})}$. If the case 2b holds as well, then $\operatorname{PoA}=\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$; otherwise, $\operatorname{PoA}=\operatorname{PoS}$.
@@
      $B_{1}<\frac{\theta B_{2}}{k+\theta^{2}}$ and $\alpha_{m-k}\geq 2\alpha_{m-k-1}$ and $\frac{2\theta}{1+\theta}\leq\frac{\alpha_{m-k}}{\alpha_{m}}\leq\frac{2(1-\theta)}{2-\theta}$ and $m-k$ is the only $2$-steep project. Then, there exists a pure strategy NE and the following holds. If the case 2a holds as well, then $\operatorname{PoS}=\frac{\alpha_{m}B_{2}+\alpha_{m-k}B_{1}}{\alpha_{m}(B_{1}+B_{2})}$; otherwise, $\operatorname{PoS}=\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$. In any case, $\operatorname{PoA}=\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$.
@@
      $B_{1}<\frac{\theta}{\left|\Omega\right|}B_{2}$ and all the project functions are equal, i.e. $\alpha_{m}=\alpha_{1}$. Then, there exist pure NE and $\operatorname{PoS}=1,\operatorname{PoA}=\frac{B_{2}}{B_{1}+B_{2}}$.
@@
    Assume that $B_{n-1}<\theta B_{n}$. Then, $\operatorname{PoA}\geq\frac{B_{n}}{\sum_{i\in\left\{1,2,\ldots,n\right\}}{B_{i}}}$. If we also have $B_{n-1}<\frac{\theta}{\left|\Omega\right|}B_{n}$ and all the project functions are equal, i.e. $\alpha_{m}=\alpha_{1}$, then there exist pure NE and $\operatorname{PoS}=1,\operatorname{PoA}=\frac{B_{n}}{\sum_{i\in\left\{1,2,\ldots,n\right\}}{B_{i}}}$.
To prove more bounds on the price of anarchy, we employ the so-called smoothness argument from Roughgarden Roughgarden (2015). This means showing that for a socially optimal profile $x^{*}$ and any profile $x$, there exist $\lambda>0$ and $\mu>-1$ such that
@@
We employ the relaxation from (Roughgarden, 2015, Remark $2.3$) that $x^{*}$ does not have to be any profile but we may pick any socially optimal one, and we use the payoff-maximization version of the argument (see (Roughgarden, 2015, Section $2.3.2$)).
@@
To summarise, we first characterise the cases of zero threshold and threshold $1$, next characterising the existence and efficiency of equilibria for $2$ players and provide some existence and efficiency results for $n\geq 2$ players. Notice that our (pure) model is not the mixing of the model where a player may invest in at most one project, because the mixing would extend the utilities linearly in the mixing coefficients (Osborne and Rubinstein, 1994, Definition 32.1), which is not the case in our (pure) model with a positive threshold.
4 Simulations
The theory provides only sufficiency results for shared effort games with more than $2$ players. To better explore the existence and efficiency of NE in these games, we simulate a variation of fictitious play Robinson (1951); intuitively, we try fictitious play because of its well-known properties of converging to an NE in some finite cases (see Section 6).
We first adapt the fictitious play Brown (1951) to our infinite game. Danskin Danskin (1981) defines the best response to maximise the average utility against all the previous strategies of the other players, while we, as well as the original fictitious play, best respond to the averaged (cumulative) strategy of the others.
Aspiring to implement the adapted fictitious play, we then suggest an algorithm for finding a best response, if it exists, for the case of two projects. For more players we prove that even best responding is already NP-hard.
4.1 Infinite-Strategy Fictitious Play for Shared Effort Games

### v2

@@ Abstract
People, robots, and companies mostly divide time and effort between projects, and shared effort games model people investing resources in public endeavors and sharing the generated values. In linear $\theta$ sharing (effort) games, a project’s value is linear in the total contribution, thus modelling predictable, uniform, and scalable activities. The threshold $\theta$ for effort defines which contributors win and receive their share, equal share modelling standard salaries, equity-minded projects, etc. Thresholds between 0 and 1 model games such as paper co-authorship and shared assignments, where a minimum positive contribution is required for sharing in the value. We constructively characterise the conditions for the existence of a pure equilibrium for $\theta\in\{0,1\}$, and for two-player games with a general threshold, and find the prices of anarchy and stability. We also provide existence and efficiency results for more than two players, and use generalised fictitious play simulations to show when a pure equilibrium exists and what its efficiency is. We propose a novel method for studying solution concepts by defining a new concept and proving its equivalence to a previously known on a large subclass of games. This means that the original concept narrows down to a more demanding concepts on certain games, providing new insights and opening a path to study both concepts conveniently. We also prove mixed equilibria always exist and bound their efficiency.
@@
Shared effort games , Contribution threshold , Equilibrium , Price of anarchy , Price of stability , fictitious play , narrowing down/broadening (a solution concept) , cyclically strong equilibrium
journal: Discrete Applied Mathematics
@@ 1 Introduction
Many real-world situations include a group of players investing resources across multiple projects. Examples of such situations include financial investments, contributions to online communities [21], Wikipedia [16], political campaigns [49], paper co-authorship [24], social exchange networks [25]. Naturally, also hobbies and attention spread fall under this category. In the formal analyses of these type of games, it is often assumed that the obtained revenue from such projects is linear in the total contribution and is shared equally [44, 24, 8]. Basically, performing standard predictable activities, which are scalable and homogeneous, often yields linear value (and we also tend to linear thinking [15]). Equal division takes place with standard salaries, fundraising for community projects, equal distribution of recognition to all the contributors, equity-minded projects (the same reward regardless the contribution), non-rivalrous values, like public parks, which can be enjoyed equally by all those who have access to them, etc. However, in most of the above examples, revenue is only shared among those who contribute at least a certain amount of effort. In this work we model this as a threshold on a player’s contribution. In particular, we analyse the situation when the considered threshold is relative to the investments of the other players. The abstract model of this paper is the first theoretical study of such shared effort games with a threshold, assuming linear project values. This provides the foundations for modelling contributions in various kinds of projects, basically parallel games, such as considering splitting time between projects, each of which is a reciprocal interaction [41].
An example with a relative threshold of $1$ is an all pay-auction. Another example is the Colonel Blotto game (see e.g. [44]), where there are several battlefields over which each of the $2$ players distributes her soldiers, and the winner at each battlefield is the one who has put more people there. A player’s payoff is the number of projects (battlefields) where she has won, the value of each project being fixed. Thus, when our model is reduced to $2$ players with a fixed value of each project, we obtain the Colonel Blotto competition, whereas we typically assume a project has a value that increases as a function of the total contribution, modelling constructive collaboration, rather than countenance. These examples are “highly thresholded” because only the player whose effort per project is maximum collects the complete revenue.
We now present examples of roughly linear growing project values, each being equally shared between all those who contribute at least a certain threshold.
  1. 1.
    Assigning points for an exercise, where a percentage of the perfect work is required to obtain the (equal) homework’s credits [20] is an absolute threshold example from a course at the University of Maryland. Linearity is a reasonable model for predictable tasks. Although our threshold is relative, while in the Maryland example it is absolute, still, professors often relate the individual grades to the average level of the students.
  2. 2.
    Consider kids selling lottery tickets for their sports club where a significant part of the income is used to improve the club’s facilities; kids that sell at least a certain number (threshold) of tickets, relatively to the others, are equally rewarded by the club with the same symbolic present (e.g., club memorabilia, like a hat, a scarf, or a jersey).
  3. 3.
    Another example is a start-up where developers are building a new tool to sell and divide the revenue equally, or any firm paying equal salaries to everyone who works enough to keep the job. Here, the resulting value is assumed to grow approximately linearly in the contributed time, when the idea being implemented is clear and no surprises are expected.
The following example with a smaller threshold is used later to further illustrate the model.
@@
It is important to understand and predict the stable contribution strategies, and suggest the efficient ones in situations where players invest resources in several projects and share the generated values. The classical stability model is the Nash equilibria, and the efficiency model of stable situations is the ratio of the least total utility of the players in an equilibrium to the optimum, called the price of anarchy ($\operatorname{PoA}$[27], and the ratio of the largest total utility in equilibrium to the optimal total utility, called the price of stability ($\operatorname{PoS}$[47, 4]. If the price of anarchy is close to $1$, then all equilibria are good, and we may follow any equilibrium profile. If the price of anarchy is low, while the price of stability is high, then we have to regulate the play by suggesting the efficient equilibria, while if even the price of stability is low, the only way to make the play socially efficient is changing the game through, for example, subsidising.
This price of anarchy for shared effort games was bounded in [8], but assuming that a player obtains at least $1/k$th of her marginal contribution, which fails to hold in a positively thresholded model. Threshold was introduced in [43], and pure equilibria for two players with close budgets and projects with linear utility functions were characterised. There is no analysis of the existence of Nash Equilibria (NE) and their efficiency in general linear shared effort games and nothing about mixed Nash equilibria is known. This paper fills this gap with the following contributions. [^4] Relatively to [43], we extend the theory also for budgets not within a threshold factor from each other, characterise the non-suppressed equilibria for threshold $1$, provide the novel concept of narrowing down Nash to the newly defined cyclically strong equilibrium, prove the existence of an NE in the mixed case, answering this natural question, and substantially extend the simulations.
Our Contributions
@@
analysing the cases of 00 threshold and $1$ threshold (Theorem 2), sufficiency results on existence of pure NE and efficiency bounds, and exploring this domain through simulation of fictitious play to find pure NE and their efficiency in 2-project multi-player games. The simulations need to be engineered based on the case in question, so the importance of the simulations is mainly in the proof of concept of implementing the infinite strategy fictitious play for 2 projects. This involves designing an $O(n\log n)$ best response algorithm for $2$-project multi-player games and proving that for more projects, best-responding is NP-hard, even for the same project functions and $2$ players.
@@
we introduce the concept of narrowing down (or broadening) a solution concept. In our case, we strengthen Nash equilibrium by adding a requirement of no profitable cyclical deviations of whole budgets, alluding to the famous top trading cycles mechanisms [1]. We then demonstrate that when no budget can make another one fall below the threshold, both solution concepts coincide.
For any number of players,
a proof of existence of mixed NE (it is an infinite game, so Nash’s Theorem [38] does not apply) and extending efficiency bounds to mixed equilibria.
After defining shared effort games in the next section, we first concentrate on pure strategies. We theoretically treat the existence and efficiency of NE for thresholds 00 and $1$ and then for games with two players in Section 3. Afterwards, we study games with any number of players. In Section 4, we define a new solution concept and narrow down Nash equilibria to that concept, whetting the appetite for defining further solution concepts and studying their relationships by analogy. Finally, we treat the mixed extension in Section 5, proving the existence of a mixed NE and showing some efficiency bounds of the pure case generalise to the mixed case. We conclude and discuss the future work in Section 7. When a proof is mostly technical, we provide its sketch and intuition in the body of the paper and defer the complete proof to A. We conduct further investigation of existence and efficiency of pure NE for more than $2$ players by employing fictitious play simulations in B. Concretely, we define an Infinite-Strategy Fictitious Play and simulate it by computing a best response, till and if it finds an equilibrium within some time. We check whether we have found an NE and if that is the case, what its efficiency is.
We introduce and study the threshold on the basic model, motivating more refinements of the model. This is a major impact, alongside the mothodological implications of the notion of narrowing down of equilibria from Section 4, which motivates defining other appropriate solution concepts and seeing when a narrowing down or a broadening occurs.
@@ 2 Model
To model investing effort in shared projects, we define shared effort games, extending the model by [8] with thresholds. The games consist of players who contribute to project, and share the value the projects generates. Henceforth, we use the general term “value” instead of the monetary “revenue”. Formally, players $N=\left\{1,\ldots,n\right\}$ contribute to projects $\Omega$. The contribution of player $i$ to project $\omega\in\Omega$ is denoted by $x^{i}_{\omega}\in{\mathbb{R}_{+}}$. Each player $i$ has a budget $B_{i}>0$, and the strategy space of player $i$ (i.e., the set of her possible actions) is [^5] We denote “defined as” by =$\stackrel{{\scriptstyle\rm\Delta}}{{=}}$. $S^{i}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{x^{i}=(x^{i}_{\omega})_{\omega\in\Omega}\in\mathbb{R}_{+}^{\left|\Omega\right|}\mid\sum_{\omega\in\Omega}{x^{i}_{\omega}}\leq B_{i}\right\}$. Assume w.l.o.g. (without loss of generality) that $B_{n}\geq\ldots\geq B_{2}\geq B_{1}>0$. Being an $\left|\Omega\right|-1$-dimensional simplex, this set is compact and convex. Denote the vector of all the contributions by $x=(x^{i}_{\omega})^{i\in N}_{\omega\in\Omega}$ and the strategies of all the players except $i$ by $x^{-i}$.
To define the utilities, each project $\omega\in\Omega$ is associated with its (really or perceivably) linear (allows concentrating on cumulative scalable cooperation across multiple “projects”, like manufacturing, technical writing, Wikipedia or environmental activism) project function, which determines its value, based on the total contribution $x_{\omega}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{i\in N}{(x_{\omega}^{i})}$ that the project receives; formally, $P_{\omega}(x_{\omega})\colon{\mathbb{R}_{+}}\to{\mathbb{R}_{+}}$, and $P_{j}(y)=\alpha_{j}y$. W.l.o.g., $\alpha_{m}\geq\alpha_{m-1}\geq\ldots\geq\alpha_{1}>0$. We denote the number of projects with the largest coefficient of project functions by $k\in\mathbb{N}$, i.e. $\alpha_{m}=\alpha_{m-1}=\ldots=\alpha_{m-k+1}>\alpha_{m-k}\geq\alpha_{m-k-1}\geq\ldots\geq\alpha_{1}$. We call those $k$ projects steep. The project’s value is distributed among the players according to the following rule. From each project $\omega\in\Omega$, each player $i$ that contributes enough gets an equal share, denoted $\phi_{\omega}^{i}(x_{\omega})\colon{\mathbb{R}_{+}}\to{\mathbb{R}_{+}}$. Such sharing is relevant to many applications where a minimum contribution is required to share the value, such as paper co-authorship and homeworks, and we study predominantly such games. For any $\theta\in[0,1]$, define the players who get a share as those who bid at least a $\theta$ fraction of the maximum bid size to $\omega$,
@@
The $\theta$-equal sharing mechanism, denoted by $M_{\text{eq}}^{\theta}$, is
$\phi^{i}_{\omega}(x_{\omega})\stackrel{{\scriptstyle\rm\Delta}}{{=}}\begin{cases}\frac{P_{\omega}(x_{\omega})}{\left|N^{\theta}_{\omega}\right|}&\mbox{if}i\in N^{\theta}_{\omega},\\ 0&\mbox{otherwise}.\\ \end{cases}$
@@
$u^{i}(x)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{\omega\in\Omega}{\phi^{i}_{\omega}(x_{\omega})}.$
@@ 3 Pure Nash Equilibrium
We begin with general results on equilibria in shared effort games. Then, we completely characterise the existence of Nash equilibria for two players, following with existence results for any number of players in Section 3.1. Finally, we provide the prices of anarchy and stability in Section 3.2. We study pure NE in this section, which are natural to play, since a pure shared effort game already has uncountably infinite strategy spaces and non-continuous utilities.
@@
A game with $M_{\text{eq}}^{0}$ admits a potential function, a pure NE exists and $\operatorname{PoA}=\operatorname{PoS}=1$.
Proof. The strategy space of player $i$ is $S^{i}$, and denote $S\stackrel{{\scriptstyle\rm\Delta}}{{=}}S^{1}\times\ldots\times S^{n}$. Define $P\colon S\to\mathbb{R}$ by $P(x)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{\omega\in\Omega}{\frac{P_{\omega}(x_{\omega})}{\left|N_{\omega}\right|}}$. This is a potential function, because it is equal to the utility of any player, and therefore, when player $i$ changes her strategy, her utility changes exactly as the potential does.
The game possesses a pure NE, whenever the potential function attains its maximum. As the linear project functions are continuous and the spaces are non-empty compact, being simplexes, so the potential function always attains its maximum (see Lemma $4.3$ in [36]).
@@
A game with $M_{\text{eq}}^{1}$ possesses NE where nobody is suppressed if and only if all the following conditions hold:
@@
If a game with $M_{\text{eq}}^{1}$ possesses NE where nobody is suppressed, then the social welfare of any such NE is $\sum_{q\leq p}{\alpha_{m_{q}}(\sum_{i:B_{i}=B_{n_{q}}}{B_{i}})}$, whereas the social optimal is $\alpha_{m_{1}}\sum_{i\in N}{B_{i}}$.
@@ 3.1 Existence of Nash Equilibrium
We are finally ready to characterise the existence of an NE for two players. The characterisation depends on the threshold, the ratio between the budgets and the highest coefficients.
@@
3.2 Efficiency
In order to facilitate decisions, it is important to analyse the efficiency of the various Nash Equilibria. We aim to find the price of anarchy ($\operatorname{PoA}$), which is the ratio of a worst NE’s efficiency to the optimum possible one, and the price of stability ($\operatorname{PoS}$), which is the ratio of a best NE’s efficiency to the optimum possible one. We first completely resolve the two-player case.
@@ 3.2 Efficiency
      $B_{1}<\frac{\theta B_{2}}{k}$ and $\frac{\alpha_{m-k}}{\alpha_{m}}\leq\min\left\{\frac{1}{1+\theta},\frac{2\theta}{1+\theta}\right\}$.
      Then, there exists a pure strategy NE and the following holds. $\operatorname{PoS}=\frac{\alpha_{m}B_{2}+\alpha_{m-k}B_{1}}{\alpha_{m}(B_{1}+B_{2})}$. If the case 2b holds as well, then $\operatorname{PoA}=\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$; otherwise, $\operatorname{PoA}=\operatorname{PoS}$.
@@
      $B_{1}<\frac{\theta B_{2}}{k+\theta^{2}}$ and $\alpha_{m-k}\geq 2\alpha_{m-k-1}$ and $\frac{2\theta}{1+\theta}\leq\frac{\alpha_{m-k}}{\alpha_{m}}\leq\frac{2(1-\theta)}{2-\theta}$ and $m-k$ is the only $2$-steep project.
      Then, there exists a pure strategy NE and the following holds. If the case 2a holds as well, then $\operatorname{PoS}=\frac{\alpha_{m}B_{2}+\alpha_{m-k}B_{1}}{\alpha_{m}(B_{1}+B_{2})}$; otherwise, $\operatorname{PoS}=\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$. In any case, $\operatorname{PoA}=\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$.
@@
      Then, there exist pure NE and $\operatorname{PoS}=1,\operatorname{PoA}=\frac{B_{2}}{B_{1}+B_{2}}$.
@@
    Assume that $B_{n-1}<\theta B_{n}$. Then, $\operatorname{PoA}\geq\frac{B_{n}}{\sum_{i\in\left\{1,2,\ldots,n\right\}}{B_{i}}}$.
    If we also have $B_{n-1}<\frac{\theta}{\left|\Omega\right|}B_{n}$ and all the project functions are equal, i.e. $\alpha_{m}=\alpha_{1}$, then there exist pure NE and $\operatorname{PoS}=1,\operatorname{PoA}=\frac{B_{n}}{\sum_{i\in\left\{1,2,\ldots,n\right\}}{B_{i}}}$.
To prove more bounds on the price of anarchy, we employ the so-called smoothness argument from Roughgarden [46]. This means showing that for a socially optimal profile $x^{*}$ and any profile $x$, there exist $\lambda>0$ and $\mu>-1$ such that
@@
We employ the relaxation from [46, Remark $2.3$] that $x^{*}$ does not have to be any profile but we may pick any socially optimal one, and we use the payoff-maximization version of the argument (see [46, Section $2.3.2$]).
@@
To summarise, we first characterise the cases of zero threshold and threshold $1$, next characterising the existence and efficiency of equilibria for $2$ players and provide some existence and efficiency results for $n\geq 2$ players. Notice that our (pure) model is not the mixing of the model where a player may invest in at most one project, because the mixing would extend the utilities linearly in the mixing coefficients [40, Definition 32.1], which is not the case in our (pure) model with a positive threshold.
4 Narrowing down Nash to Cyclically Strong Equilibrium
We now introduce a new generic method to study solution concepts. Having introduced a new solution concept by relaxing or strengthening some properties of an existing solution concept, imagine it subsequently turns out that both concepts coincide on a subclass of games. Now, if the new concept generally strengthens the previous one, then we conclude that on the subclass where both concepts are equivalent, the requirements of the original concept imply the new concept, implying a narrowing down of the original concept to the new one. On the other hand, if the new concept relaxes the original one, then their equivalence on a certain subset of games means that already the lighter requirements imply the original concept, demonstrating a broadening of the original concept. This also allows characterising one solution using the equivalent one, and to transfer the efficiency bounds.
Here, we are going to (apparently) strengthen the Nash equilibria on thresholded shared effort games by adding the requirement that no cyclical deviation of whole budgets is allowed. Cyclical moves are practically and theoretically important, such as the famous top trading cycles mechanisms [1]. Subsequently, we will prove that the resulting solution concept is actually equivalent to Nash equilibrium on thresholded shared effort games where no budget can suppress another one, thereby constituting a narrowing down of Nash equilibrium. In other words, having no profitable unilateral deviations already implies possessing no cyclical deviation of whole budgets. This exemplifies a new useful technique we hope will serve many researchers in the future.
Formally, given a shared effort game, we define